OpenStudy (anonymous):

find all the critical point of f(x) = 3x^(5/3) - 15x^(2/3)

terenzreignz (terenzreignz):

Critical points are points where the derivative is zero.

terenzreignz (terenzreignz):

Of course, differentiate first :)

OpenStudy (anonymous):

ok.. i got x=0, x=2

terenzreignz (terenzreignz):

let's see.... \[\Large \frac{d}{dx}\left(3x^{\frac53}-15x^{\frac23}\right)=5x^{\frac23}-10x^{-\frac13}\]

OpenStudy (anonymous):

is my answer correct?

terenzreignz (terenzreignz):

Try 0. the derivative doesn't exist, because you have... \[\Large 5x^{\frac23}-10x^{-\frac13}=5x^{\frac23}-\frac1{10x^{\frac13}}\] an x in the denmoinator.

OpenStudy (anonymous):

so that will be disregarded right? so i will have 2 for the value of x

terenzreignz (terenzreignz):

so yes, that is a critical point (sorry, I forgot to mention that points where the derivative does not exist are also critical points)

OpenStudy (anonymous):

owh.. i see

OpenStudy (anonymous):

so what next? how will i graph?

terenzreignz (terenzreignz):

Hang on, I'm not sure about the 2 yet. Let's try equating the derivative to zero. \[\Large0=5x^{\frac23}-\frac1{10x^{\frac13}}\]\[\Large5x^{\frac23}=\frac1{10x^{\frac13}}\]\[\Large5x^{\frac23}\cdot 10x^{\frac13}=1\]\[\Large50x=1\] Are you sure it's x = 2? :P

OpenStudy (anonymous):

can you factor out 5x^(1/3) from 5x^(2/3)−10x^(−1/3)?

terenzreignz (terenzreignz):

You can... but why would you do that?

OpenStudy (anonymous):

what will be the answer if you would factor out that one? 5x^(-1/3) i got a little confused

terenzreignz (terenzreignz):

I think it leads nowhere :)

OpenStudy (anonymous):

owh.. lol.. so x = 1/50?

terenzreignz (terenzreignz):

yeah :)

OpenStudy (anonymous):

what is the answer if i would take out 5x^(-1/3) from 5x^(2/3) i really am confused on what will be the outcome

terenzreignz (terenzreignz):

Hang on, I may have made a mistake.

terenzreignz (terenzreignz):

Okay... apparently, 2 was right.

terenzreignz (terenzreignz):

sorry.

OpenStudy (anonymous):

yehey :D so.. what to do next ?

terenzreignz (terenzreignz):

just like that? You're not going to ask me where my error was? :/ You didn't spot it yet... means you may well commit the same... come on, a little challenge :P

terenzreignz (terenzreignz):

\[\Large0=5x^{\frac23}-\frac1{10x^{\frac13}}\] I should not have put 10 in the denominator. Big mistake. Correct would be \[\Large0=5x^{\frac23}-\frac{10}{x^{\frac13}}\color{green}\checkmark \]

OpenStudy (anonymous):

you shouldn't include 10 in the denominator.. it should be on the numerator :D

terenzreignz (terenzreignz):

yeah that :) sorry about that. I might be getting drowsy.

terenzreignz (terenzreignz):

And as far as I know... we're done.

OpenStudy (anonymous):

that's ok.. :))

terenzreignz (terenzreignz):

We already have the critical points.

OpenStudy (anonymous):

then how to graph?

terenzreignz (terenzreignz):

whoops... you'll need to consult one of the big guys for that :3 I'm rather terrible at graphing, you see...

OpenStudy (anonymous):

owh.. that'll be okay then.. thanks for the help! :D

terenzreignz (terenzreignz):

No problem :)